OpenStudy (anonymous):

Can someone help me solve.. ? two times the cube root of x-1 to the fourth plus four is equal to 36. I have to solve for x. 2*3√(x-1)^4 +4 = 36

OpenStudy (hba):

@TRowstudy Welcome to OS.

OpenStudy (hba):

\[ 2 \times \sqrt[3]{(x-1)^4}+4=36\]

OpenStudy (hba):

1)Subtract 4 from both sides

OpenStudy (anonymous):

Oh thank you.

OpenStudy (hba):

What do you get when you do that ?